OpenStudy (surryyy):

Find the measure of angle A in the triangle below. Answers: 14 degrees 28 degrees 52 degrees 100 degrees

OpenStudy (surryyy):



OpenStudy (midhun.madhu1987):

The sum of the 3 angles of a triangle is 180

OpenStudy (midhun.madhu1987):

So, A + B + C = 180 2x + 4x - 4 + 100 = 180 Find x.

OpenStudy (surryyy):

x = 14

OpenStudy (midhun.madhu1987):

Correct :)

OpenStudy (surryyy):

Thnx.

Directrix (directrix):

Find the measure of angle A in the triangle below. x = 14 but you are asked to find Angle A.

Directrix (directrix):

m<A = 2*x = 2*14 = ?

Directrix (directrix):

@midhun.madhu1987 Please check answer. Thanks.

OpenStudy (midhun.madhu1987):

yeah... 2x = 2 * 14 =

OpenStudy (midhun.madhu1987):